find y so the distance between (2,y) and (1,2) is sqrt 2
Answer by lwsshak3(11628)

find y so the distance between (2,y) and (1,2) is sqrt 2
distance formula: d=√((y1-y2)^2+(x1-x2)^2)
√2=√(y-2)^2+(2-1)^2
√2=√(y^2-4y+4+1)=√(y^2-4y+5)
square both sides:
2=y^2-4y+5
y^2-4y+3=0
(y-3)(y-1)=0
y=3
or
y=1